OpenStudy (anonymous):

I need these: One function, f(x), with two real rational solutions. One function, g(x), with two real irrational solutions. One function, h(x), with two complex solutions.

OpenStudy (anonymous):

@SolomonZelman @issy987 @ranga

OpenStudy (watchmath):

Does f(x),g(x),h(x) needs to be polynomia with integer coefficients?

OpenStudy (anonymous):

No, they do not

OpenStudy (ranga):

I will do the first one: f(x) = (x - 2)(x - 3) = x^2 - 5x + 6 has two real solutions

OpenStudy (anonymous):

Okay, but why does it have two real solutions?

OpenStudy (ranga):

Solution means setting f(x) to zero and solving for x. (x - 2)(x - 3) = 0 will be 0 when x =2 or x = 3. 2 and 3 are real solutions.

OpenStudy (anonymous):

Okay, that you @ranga for explaining that.

OpenStudy (ranga):

The general form of quadratic equation is ax^2 + bx + c = 0 Do you know what the discriminant is for a quadratic expression?

OpenStudy (anonymous):

No, I do not. My teacher hasn't really been much of a help.

OpenStudy (ranga):

\[\Large ax ^{2} + bx + c = 0\] Let D be Discriminant. Then, \[\Large D = b ^{2} - 4ac\] The roots or solutions of a quadratic equation are: \[\Large \frac{ -b \pm \sqrt{D} }{ 2a }\] If D is positive, the equation will have two real roots If D is zero, the equation will have two equal roots If D is negative, the equation will have two complex roots

OpenStudy (anonymous):

Okay, wow, this is a whole lot better learning then how my teacher put it. Are you a teacher?

OpenStudy (ranga):

No.

OpenStudy (anonymous):

Should be :)

OpenStudy (ranga):

The first example I gave you has 2 and 3 as solutions. 2 and 3 are real and they are rational numbers. For the second function g(x), they want two irrational solutions. Numbers like: \[\Large \pi, \sqrt{2}, \sqrt{7}\] are irrational numbers. They are still real numbers but they are irrational because they cannot be put in the form of a fraction such m/n. So to have two solutions that are irrational we can multiply: \[\Large g(x) = (x - \sqrt{2})(x - \sqrt{5}) = x ^{2} - (\sqrt{5} + \sqrt{2})x + \sqrt{10}\]

OpenStudy (anonymous):

Okay, this is actually starting to make some sense to me now.

OpenStudy (anonymous):

\[x=-3+\sqrt{(3)^2-4(3)}\div2\]

OpenStudy (anonymous):

So would that be a complex solution?

OpenStudy (ranga):

For two complex solutions choose a, b and c of the quadratic equation such that the discriminant D is negative. The solution to the quadratic equation given earlier requires taking the square root of D. If D is negative you cannot a real square root of D and so the solutions will be complex.

OpenStudy (ranga):

Yes. Except the divided by 2 should be below the entire thing not just the square root of the discriminant. And also, they want the function h(x). You possibly chose a = 1, b = 3 and c = 3 so h(x) = ax^2 + 3x + 3 will have two complex roots.

OpenStudy (anonymous):

Yes, I didnt know how to put the 2 under the equation.

OpenStudy (ranga):

And the + in your solution should be + or -

OpenStudy (anonymous):

Okay, thank you so much for your help :) Learned more from you in the past half hour than I do two hours with my teacher

OpenStudy (ranga):

You are welcome. Glad to be of some assistance.

OpenStudy (ranga):

Oh, two other subsets of info that would be useful under the case the discriminant is positive: If the discriminant is a perfect square there will be two rational real roots If the discriminant is not a perfect square there will be two irrational real roots

OpenStudy (anonymous):

Okay, so how would I convert something like f(x) = (x - 2)(x - 3) = x^2 - 5x + 6 from standard form to vertex form?

OpenStudy (ranga):

You take x^2 - 5x + 6 and complete the square. Are you familiar with it?

OpenStudy (anonymous):

What square are you referring to?

OpenStudy (ranga):

Taking ax^2 + bx + c and expressing it in the form a(x - d)^2 + e is usually referred to as "completing the square"

OpenStudy (anonymous):

Okay, so I would plug that into the f(x) equation?

OpenStudy (ranga):

You will take the f(x) expression in the standard form and try to express it in the vertex form by completing the square. x^2 - 5x + 6 = (x - 5/2)^2 - 25/4 + 6 Do you follow this step?

OpenStudy (anonymous):

I believe so

OpenStudy (ranga):

That is all there is to it. Can you try to put this in vertex form: 3x^2 + 2x - 1 ?

OpenStudy (anonymous):

I will try

OpenStudy (ranga):

Factor the 3 out of the entire expression first. It will make it easier to work with.

OpenStudy (ranga):

3x^2 + 2x - 1 = 3(x^2 + 2/3x - 1/3)

OpenStudy (anonymous):

Okay, so is the answer what you just wrote or will the answer come from that?

OpenStudy (ranga):

By factoring out the 3 you have made a to be 1 in the quadratic expression. a = 1, b = 2/3, c = -1/3 Once you have made a to be 1, just divide b by 2 to get the term inside the square here b/2 = (2/3) / 2 = 1/3 so x^2 + 2/3x - 1/3 = (x - 1/3)^2 - 1/9 - 1/3 Follow?

OpenStudy (anonymous):

Okay, written like that I now see what is supposed to be done

OpenStudy (ranga):

so x^2 + 2/3x - 1/3 = (x - 1/3)^2 - 1/9 - 1/3 = (x - 1/3)^2 - 4/9 So the original function 3x^2 + 2x - 1 = 3[ (x - 1/3)^2 - 4/9 ] The reason you take a quadratic function such as ax^2 + bx + c and put it in the vertex form: a(x - h)^2 + k is: If you graph the function f(x) = ax^2 + bx + c you will see it is a parabola. One of the frequently asked questions is what is the coordinates of the vertex of the parabola. So if you put ax^2 + bx + c in the form a(x - h)^2 + k then the vertex will be (h, k).

OpenStudy (anonymous):

Okay, so first step in putting f(x) = (x - 2)(x - 3) = x^2 - 5x + 6 into vertex form would be to put x^2 - 5x + 6 in to complete the square

OpenStudy (ranga):

Yes. Let us take this problem: Find the vertex of the parabola f(x) = 5x^2 + 4x - 1 The first step is factor the 5 out to make a = 1 in the quadratic expression. 5x^2 + 4x - 1 = 5(x^2 + 4/5x - 1/5) So b here is 4/5. Divide it by 2 to get 2/5. This 2/5 will go as -2/5 inside the square:

OpenStudy (ranga):

5(x^2 + 4/5x - 1/5) = 5[ (x - 2/5)^2 - 4/25 - 1/5]

OpenStudy (anonymous):

Okay, So after the\at, what would you do?

OpenStudy (ranga):

Simplify the expression on the right: 5[ (x - 2/5)^2 - 4/25 - 1/5 ] = 5[ (x - 2/5)^2 - 9/25 ] Take the factor 5 inside (remember, we want the function to look like a(x-h)^2 + k: 5[ (x - 2/5)^2 - 9/25 ] = 5(x - 2/5)^2 - 9/5 h = 2/5 and k = -9/5 So the vertex is: (2/5, -9/5)

OpenStudy (anonymous):

Wow, this is defiantly going to pay off on my quiz tomorrow.

OpenStudy (ranga):

Good luck! I have to go now.

OpenStudy (anonymous):

Well, I am absolutely grateful for your help :)

OpenStudy (ranga):

I just noticed there was a sign mistake (+/-) in my earlier answer which I am correcting here. Let us take this problem: Find the vertex of the parabola f(x) = 5x^2 + 4x - 1 The first step is factor the 5 out to make a = 1 in the quadratic expression. 5x^2 + 4x - 1 = 5(x^2 + 4/5x - 1/5) So b here is 4/5. Divide it by 2 to get 2/5. This 2/5 will go as +2/5 inside the square. If the coefficient of the x term is positive, the square term will be of the form: (x + m)^2 If the coefficient of the x term is negative, the square term will be of the form: (x - m)^2 5(x^2 + 4/5x - 1/5) = 5[ (x + 2/5)^2 - 4/25 - 1/5 ] = 5[ (x + 2/5)^2 - 9/25 ] Take the 5 inside: 5(x + 2/5)^2 - 9/5 Compare it to a(x - h)^2 + k h = -2/5 and k = -9/5 So the vertex is: (-2/5, -9/5)

OpenStudy (ranga):

Find the vertex of 3x^2 + 2x - 1 3x^2 + 2x - 1 = 3(x^2 + 2/3x - 1/3) = 3[ (x + 1/3)^2 - 1/9 - 1/3 ] = 3[ (x + 1/3)^2 - 4/9 ] = 3(x + 1/3)^2 - 4/3 Compare it to a(x - h)^2 + k h = -1/3 and k = -4/3. Vertex = (-1/3, -4/3)